### v2

# Square-free powers of Cohen-Macaulay simplicial forests

Abstract.
Let $I(\Delta)^{[k]}$ denote the $k^{\text{th}}$ square-free power of the facet ideal of a simplicial complex $\Delta$ in a polynomial ring $R$. Square-free powers are intimately related to the ‘Matching Theory’ and ‘Ordinary Powers’. In this article, we show that if $\Delta$ is a Cohen-Macaulay simplicial forest, then $R/I(\Delta)^{[k]}$ is Cohen-Macaulay for all $k\geq 1$. This result is quite interesting since all ordinary powers of a graded radical ideal can never be Cohen-Macaulay unless it is a complete intersection. To prove the result, we introduce a new combinatorial notion called special leaf, and using this, we provide an explicit combinatorial formula of $\mathrm{depth}(R/I(\Delta)^{[k]})$ for all $k\geq 1$, where $\Delta$ is a Cohen-Macaulay simplicial forest. As an application, we show that the normalized depth function of a Cohen-Macaulay simplicial forest is nonincreasing.
Key words and phrases:
Square-free powers, Cohen-Macaulay rings, simplicial forests, depth
2020 Mathematics Subject Classification:
Primary: 13H10, 13C15; Secondary: 05E40, 13F55
1. Introduction
A classical way to treat the class of square-free monomial ideals is by identifying them as a Stanley-Reisner ideal of simplicial complexes. This approach has been instrumental in bridging several areas of mathematics and produced many fascinating results. It was Faridi [10] who first treated square-free monomial ideals as a facet ideal of simplicial complexes and introduced the notion of simplicial trees. In case of simple graphs, the class of trees is nothing but the one-dimensional simplicial trees.
Over the years, many algebraic properties of edge ideals of trees have been extended to facet ideals of simplicial trees. For instance, Faridi [10] generalized the sliding depth property of edge ideals of trees proved by Simis-Vasconcelos-Villarreal [21] to the facet ideals of simplicial trees. Moreover, these ideals have normal and Cohen-Macaulay Rees  rings.
A central theme in combinatorial commutative algebra is understanding how properties of graded ideals extend to their powers. Many algebraic invariants exhibit predictable asymptotic behavior. A classical example is the work of Herzog, Hibi, and Zheng [15], who extended the linearity of minimal free resolutions of edge ideals of simple graphs to all of their ordinary powers. If we consider the Cohen-Macaulay property of a square-free monomial ideal, it is known that all of its ordinary powers are Cohen-Macaulay only when the ideal is a complete intersection ideal. This situation becomes much more intricate if we consider other kind of powers, such as symbolic powers [18], or square-free powers [13]. This article primarily focuses on the square-free powers of facet ideals of simplicial forests.
Let $I\subseteq R=\mathbb{K}[x_{1},\ldots,x_{n}]$ be a square-free monomial ideal, where $\mathbb{K}$ is a field. The concept of square-free powers was introduced in [1] and later extended to all monomial ideals through matching powers [6]. The motivation for studying the square-free powers comes from graph theory. The $k^{th}$ square-free power of $I$, denoted by $I^{[k]}$, is the ideal of $R$ generated by all square-free monomials of the ideal $I^{k}$. Note that $I^{[k]}=\langle 0\rangle$ for $k>>0$. Let $\mathcal{G}(I)$ denote the unique minimal monomial generating set of a monomial ideal $I$. Then one can observe that $\mathcal{G}(I^{[k]})$ is the set of square-free monomials contained in $\mathcal{G}(I^{k})$. In other words, the minimal generating set $\mathcal{G}(I^{[k]})$ of the $k^{th}$ square-free power of $I$ has a one-to-one correspondence with the set of all possible $k$-matchings of the underlying simplicial complex. So far, numerous studies have been conducted in various directions on the square-free powers of edge ideals of graphs (see, for instance, [4, 5, 7, 8, 9, 12, 13, 20]). However, to the best of our knowledge, the investigation of square-free powers of arbitrary square-free monomial ideals has been explored only in [16].
The primary goal of this article is to explore the Cohen-Macaulay property of square-free powers of facet ideals of simplicial forests. It is well-known that for any graded radical ideal $I$ in a polynomial ring $R$, $R/I^{k}$ is Cohen-Macaulay for all $k$ if and only if $I$ is a complete intersection (see, for instance, [3]). This naturally raises the question of whether the same holds true for square-free powers. Interestingly, this is not the case, as it has been shown in [5] that all square-free powers of the edge ideals of Cohen-Macaulay forests are also Cohen-Macaulay. Meanwhile, in [13], Ficarra and Moradi showed that among the classes of chordal, Cameron-Walker, and very well-covered graphs, the only graphs whose all square-free powers of edge ideals satisfy the Cohen-Macaulay property are Cohen-Macaulay forests and complete graphs. To the best of our knowledge, no such square-free monomial ideals, apart from the above-mentioned edge ideals of simple graphs, are known for which all square-free powers are Cohen-Macaulay. Identifying these classes of ideals remains a challenging problem. In this paper, we provide a broad class of such square-free monomial ideals, which are not necessarily equigenerated. In particular, as a main result of this paper, we show that all square-free powers of the facet ideal of a Cohen-Macaulay simplicial forest are Cohen-Macaulay.
To achieve our main result, we explicitly compute the (Krull) dimension and depth of $R/I(\Delta)^{[k]}$ for a Cohen-Macaulay simplicial forest $\Delta$. In this context, we introduce the notion of special leaves in a simplicial complex and show their existence in Cohen-Macaulay simplicial forests. Here, we remark that the concept of special leaves might be useful in further studies of simplicial forests. The computation of $\mathrm{depth}(R/I(\Delta)^{[k]})$ intricately uses the combinatorial characterization of Cohen-Macaulay simplicial forests and the presence of a special leaf, along with the standard homological techniques. The numerical formula of the depth of $R/I(\Delta)^{[k]}$ also helps us to analyze the behavior of the normalized depth function. The normalized depth function of square-free powers was introduced by Erey-Herzog-Hibi-Madani in [8], where they conjectured that this function is nonincreasing. Later, Fakhari disproved this conjecture in [19]. However, the conjecture still remains open for square-free powers of edge ideals of graphs and holds affirmatively for several subclasses (see, for instance, [4, 13]). It follows from our numerical formula for the depth that the conjecture holds affirmatively in the case of Cohen-Macaulay simplicial forests.
The paper is structured as follows. In Section˜2, we recall the definitions and basic properties of a Cohen-Macaulay simplicial forest. The primary objective of this section is to introduce the concept of a special leaf in a simplicial complex. Further, we show the existence of such a special leaf in a Cohen-Macaulay simplicial forest (see Lemma˜2.10). In Section˜3, we first prove two technical lemmas Lemma˜3.1 and Lemma˜3.4. Then, we establish the formula for the Krull dimension of the corresponding quotient rings of the square-free powers of a Cohen-Macaulay simplicial forest (see Theorem˜3.5). Finally, we derive the depth formula as follows.
Theorem 3.6.
Let $\Delta$ be a Cohen-Macaulay simplicial forest on the vertex set $V(\Delta)$. Then
$\mathrm{depth}(R/I(\Delta)^{[k]})=|V(\Delta)|-\nu(\Delta)+k-1,$
where $\nu(\Delta)$ is the matching number of $\Delta$.
As an immediate application of the above theorem and the computation of dimension, we obtain the main result of the paper:
Corollary 3.7.
Let $\Delta$ be a Cohen-Macaulay simplicial forest. Then for all $k\geq 1$, $I(\Delta)^{[k]}$ is also Cohen-Macaulay.
Additionally, in Theorem˜3.10, we show that the normalized depth function of square-free powers of the facet ideal of a Cohen-Macaulay simplicial forest is nonincreasing.
2. Cohen-Macaulay Simplicial Forests and Special Leaf
In this section, we begin by reviewing the definition of simplicial forests and the combinatorial criteria for their Cohen-Macaulay property. The main purpose of this section is to introduce a new combinatorial concept in a simplicial complex called special leaf. Also, we show the existence of such a special leaf in a Cohen-Macaulay simplicial forest, and this fact is one of the key ingredients in proving our main result in the subsequent section.
A simplicial complex $\Delta$ over a set of vertices $V=\{x_{1},\ldots,x_{n}\}$ is a collection of subsets of $V$ such that if $F\in\Delta$ and $G\subseteq F$, then $G\in\Delta$. An element of $\Delta$ is called a face of $\Delta$, and the dimension of a face $F$ of $\Delta$, denoted by $\mathrm{dim}(F)$, is defined as |F|1|F|-1, where $|F|$ is the number of vertices in $F$. The dimension of $\Delta$, denoted by $\dim(\Delta)$, is given by $\mathrm{max}\{\mathrm{dim}(F)\mid F\in\Delta\}$. The maximal faces of $\Delta$ under inclusion are called the facets of $\Delta$. We denote the set of all facets of $\Delta$ by $\mathcal{F}(\Delta)$. If $\mathcal{F}(\Delta)=\{F_{1},\ldots,F_{m}\}$, then we simply write $\Delta=\langle F_{1},\ldots,F_{m}\rangle$. For each $F_{i}\in\mathcal{F}(\Delta)$, the set $\mathcal{N}_{\Delta}(F_{i})=\{F_{j}\in\mathcal{F}(\Delta)\mid F_{j}\cap F_{i}\neq\emptyset\text{ and }j\neq i\}$ is called the set of neighbors of $F_{i}$. The set $\mathcal{N}_{\Delta}[F_{i}]=\mathcal{N}_{\Delta}(F_{i})\cup\{F_{i}\}$ is called the closed neighborhood of $F_{i}$ in $\Delta$. In this article, by a subcomplex of $\Delta$, we mean a simplicial complex $\Delta^{\prime}$ where $\mathcal{F}(\Delta^{\prime})\subseteq\mathcal{F}(\Delta)$ following [10]. Note that this differs from the standard notion of subcomplex commonly used in the literature.
2.1. Simplicial forests and their facet ideals
In [10], Faridi extended the notion of forests in graph theory to the realm of simplicial complexes and introduced the class of simplicial forests. A facet $F$ of a simplicial complex $\Delta$ is called a leaf if either $F$ is the only facet of $\Delta$ or there exists a facet $G\in\Delta$ such that $G\neq F$ and $F\cap G\supseteq F\cap H$ for all facets $H\neq F$ of $\Delta$. Such a facet $G$ is then called a joint of $F$ in $\Delta$. A vertex of a simplicial complex $\Delta$ is called a free vertex if it belongs to exactly one facet of $\Delta$. Note that every leaf $F$ of $\Delta$ contains a free vertex. A connected simplicial complex $\Delta$ is a simplicial tree if every nonempty subcomplex of $\Delta$ has a leaf. A simplicial forest is a simplicial complex whose every connected component is a simplicial tree. It directly follows from the definition that if $\Delta$ is a simplicial tree, then any subcomplex of $\Delta$ is a simplicial forest.
Let $\Delta$ be a simplicial complex with the set of facets $\mathcal{F}(\Delta)$. A subset $M$ of $\mathcal{F}(\Delta)$ is called a matching of $\Delta$ if for any two distinct facets $F,F^{\prime}\in M$, we have $F\cap F^{\prime}=\emptyset$. The maximum cardinality of a set of matchings in $\Delta$ is called the matching number of $\Delta$ and is denoted by $\nu(\Delta)$. Let $V(\Delta)=\{x_{1},\ldots,x_{n}\}$, and $R$ denote the polynomial ring $\mathbb{K}[x_{1},\ldots,x_{n}]$, where $\mathbb{K}$ is a field. Sometimes $R$ is also denoted by $\mathbb{K}[V(\Delta)]$ in order to specify the relationship with the vertex set $V(\Delta)$. For a subset of variables $S\subseteq\{x_{1},\ldots,x_{n}\}$, we will write $\mathbf{x}_{S}$ to denote the monomial $\prod_{x_{i}\in S}x_{i}$. Then the facet ideal of $\Delta$, denoted by $I(\Delta)$, is the ideal generated by monomials corresponding to each facet of $\Delta$, i.e.,
$I(\Delta)=\langle\mathbf{x}_{F}\mid F\text{ is a facet of }\Delta\rangle.$
2.2. Cohen-Macaulay simplicial forests
The Cohen-Macaulay property of a simplicial forest is combinatorially characterized in terms of the notion of grafting in [11]. In fact, the procedure called grafting of a simplicial complex was first introduced by Faridi [11], which we describe below.
Definition 2.1.
[11, Definition 7.1] A simplicial complex $\Delta$ is said to be a grafting of a simplicial complex $\Delta^{\prime}=\langle G_{1},\ldots,G_{s}\rangle$ with the simplices $F_{1},\ldots,F_{r}$ (or we say that $\Delta$ is grafted) if $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ with the following properties:
  1. (1)
    $V(\Delta^{\prime})\subseteq\cup_{i=1}^{r}F_{i}$,
  2. (2)
    $F_{1},\ldots,F_{r}$ are all the leaves of $\Delta$,
  3. (3)
    $\{F_{1},\ldots,F_{r}\}\cap\{G_{1},\ldots,G_{s}\}=\emptyset$,
  4. (4)
    For $i\neq j$, $F_{i}\cap F_{j}=\emptyset$,
  5. (5)
    If $G_{i}$ is a joint of $\Delta$, then $\Delta\setminus\langle G_{i}\rangle$ is also grafted. Equivalently, for each $G_{i}\in\mathcal{F}(\Delta^{\prime})$, $\Delta\setminus\langle G_{i}\rangle$ is a grafted simplicial complex.
One of the main results in this direction is the following combinatorial classification of the Cohen-Macaulay property of the facet ideal of a simplicial tree.
Theorem 2.2.
[11, Theorem 8.2, Corollary 7.8] Let $\Delta$ be a simplicial tree, and let $I(\Delta)$ denote its facet ideal in the polynomial ring $R$. Then $R/I(\Delta)$ is Cohen-Macaulay if and only if $\Delta$ is grafted.
Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ be a grafting of the simplicial complex $\Delta^{\prime}=\langle G_{1},\ldots,G_{s}\rangle$. If for some $j\in[s]$ and $i\in[r]$, $G_{j}\subseteq F_{i}$, then we see that $\Delta$ can also be considered as a grafting of the simplicial complex $\Delta^{\prime\prime}=\langle G_{1},\ldots,\widehat{G_{j}},\ldots,G_{s}\rangle$. Thus, without loss of generality, we may assume that $F_{1},\ldots,F_{r},G_{1},\ldots,G_{s}$ are all the facets of $\Delta$, i.e., $\mathcal{F}(\Delta)=\{F_{1},\ldots,F_{r},G_{1},\ldots,G_{s}\}$. Such a presentation of $\Delta$ is said to be a minimal presentation, and it is unique up to a permutation of the indices.
Remark 2.3.
From now onwards, when we say that $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ is a Cohen-Macaulay simplicial forest, we assume that $\Delta$ has the minimal presentation and $\Delta$ is a grafting of the simplicial forest $\Delta^{\prime}=\langle G_{1},\ldots,G_{s}\rangle$ by the simplices $F_{1},\ldots,F_{r}$.
The notion of a good leaf was introduced in [2]. However, the existence of such a leaf in an arbitrary simplicial forest was first proved in [14].
Definition 2.4 (Good leaf).
A facet $F$ of a simplicial complex $\Delta$ is called a good leaf if $F$ is a leaf of every subcomplex of $\Delta$ that contains the facet $F$.
The equivalent definition of a good leaf given below is well-known in the literature, but we include a short proof here for the sake of completeness.
Proposition 2.5.
Let $\Delta$ be a simplicial complex and $F\in\Delta$ be a leaf with $\mathcal{N}_{\Delta}(F)=\{F_{1},\ldots,F_{m}\}$. Then $F$ is a good leaf of $\Delta$ if and only if there is a permutation $\{j_{1},\ldots,j_{m}\}$ of the set $\{1,\ldots,m\}$ such that $F\cap F_{j_{1}}\supseteq F\cap F_{j_{2}}\supseteq\dots\supseteq F\cap F_{j_{m}}$.
Proof.
The ‘only if’ part of the proof follows from the definition of good leaf. Conversely, without loss of generality, assume that $F\cap F_{1}\supseteq F\cap F_{2}\supseteq\dots\supseteq F\cap F_{m}$. Let $\Delta^{\prime}$ be a subcomplex of $\Delta$ containing $F$. If $F\cap G=\emptyset$ for each $F\neq G\in\mathcal{F}(\Delta^{\prime})$, then $F$ is clearly a leaf of $\Delta^{\prime}$. Otherwise, take $l=\mathrm{min}\{i\mid F_{i}\in\Delta^{\prime}\}$. Then we have $F\cap F_{l}\supseteq F\cap G$ for all $G\in\mathcal{F}(\Delta^{\prime})$. Therefore, $F$ is a leaf in $\Delta^{\prime}$. ∎
For a Cohen-Macaulay simplicial forest, it follows from Theorem˜2.2 that each leaf is a good leaf.
Lemma 2.6.
[11] Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ be a Cohen-Macaulay simplicial forest. Then for each $1\leq i\leq r$, $F_{i}$ is a good leaf of $\Delta$.
Remark 2.7.
For a Cohen-Macaulay simplicial forest $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$, we have $\nu(\Delta)=r$. Indeed, since $F_{i}\cap F_{j}=\emptyset$ for each $i\neq j$, we have $\nu(\Delta)\geq r$. On the other hand, if $\{H_{1},\ldots,H_{m}\}$ is a matching of $\Delta$ such that $\nu(\Delta)=m$, then by (1) in Definition˜2.1, for each $i\in[m]$, there exists some $j_{i}\in[r]$ such that $H_{i}\cap F_{j_{i}}\neq\emptyset$. Our aim is to show that $F_{j_{i}}\neq F_{j_{l}}$ for each $i\neq l$. On the contrary, suppose $F_{j_{i}}=F_{j_{l}}$. Since $H_{i}\cap H_{l}=\emptyset$, we have $H_{i},H_{l}\in\{G_{1},\ldots,G_{s}\}$. Therefore, by Proposition˜2.5 and Lemma˜2.6, we have $F_{j_{i}}\cap H_{i}\subseteq F_{j_{l}}\cap H_{i}$ or $F_{j_{l}}\cap H_{i}\subseteq F_{j_{i}}\cap H_{i}$, a contradiction to the fact that $H_{i}\cap H_{j}=\emptyset$. Thus $m\leq r$ and we have $\nu(\Delta)=r$.
2.3. Special leaf
While the presence of a good leaf in any simplicial forest is helpful in this context, calculating the depth of square-free powers of facet ideals of Cohen-Macaulay simplicial forests explicitly requires the introduction of a new concept called special leaf.
Definition 2.8 (Special leaf).
Let $\Delta$ be a simplicial complex with the set of facets $\mathcal{F}(\Delta)$. A leaf $F$ of $\Delta$ is said to be a special leaf if the following holds:
[…]
where $H,H^{\prime}\in\mathcal{F}(\Delta)\setminus\{F\}$.
[…]
$x_{1}$$x_{2}$$x_{3}$$y_{3}$$y_{4}$$y_{1}$$y_{2}$$x_{4}$$\Delta_{2}$
Figure 1. Simplicial trees and special leaf
Unlike a good leaf, a simplicial forest may not contain any special leaf. For instance, no leaf in the simplicial forest $\Delta_{1}$ in Figure˜1 is a special leaf. Indeed, if we set $F_{1}=\{x,y_{1},y_{2}\},F_{2}=\{x,y_{3},y_{4}\}$ and $F_{3}=\{x,y_{5},y_{6}\}$, then $(F_{1}\cap F_{2})\setminus F_{3}=\emptyset$, but $F_{1}\cap F_{2}\neq\emptyset$ in $\Delta_{1}$. Similarly, $(F_{1}\cap F_{3})\setminus F_{2}=\emptyset$, but $F_{1}\cap F_{3}\neq\emptyset$, and $(F_{2}\cap F_{3})\setminus F_{1}=\emptyset$, but $F_{2}\cap F_{3}\neq\emptyset$ in $\Delta_{1}$. Now, consider the simplicial complex $\Delta_{2}$ in Figure˜1, which is a Cohen-Macaulay simplicial tree. Note that the leaves $\{x_{1},y_{1}\},\{x_{2},y_{2}\}$ and $\{x_{4},y_{4}\}$ are special leaves. However, the leaf $\{x_{3},y_{3}\}$ in $\Delta_{2}$ is not a special leaf since $\{x_{3},x_{4}\}\cap\{x_{1},x_{2},x_{3}\}\neq\emptyset$ but $(\{x_{3},x_{4}\}\cap\{x_{1},x_{2},x_{3}\})\setminus\{x_{3},y_{3}\}=\emptyset$. Thus, not every leaf in a Cohen-Macaulay simplicial forest is a special leaf.
Remark 2.9.
Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ be a Cohen-Macaulay simplicial forest with the set of leaves $\{F_{1},\ldots,F_{r}\}$. Since $F_{i}\cap F_{j}=\emptyset$, it is easy to see that a leaf $F$ of $\Delta$ is a special leaf if and only if the following implication holds:
$G_{l}\cap G_{m}\neq\emptyset\ \Rightarrow\ (G_{l}\cap G_{m})\setminus F\neq\emptyset,$
where $l,m\in[s]$ and $l\neq m$.
We conclude this section by showing the existence of a special leaf in a Cohen-Macaulay simplicial forest.
Lemma 2.10.
Let $\Delta$ be a Cohen-Macaulay simplicial forest. Then, $\Delta$ contains at least one special leaf.
Proof.
Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ be a Cohen-Macaulay simplicial forest, and set $\Delta^{\prime}=\langle G_{1},\ldots,G_{s}\rangle$. Take any good leaf of $\Delta^{\prime}$, say $G_{1}\in\mathcal{F}(\Delta^{\prime})$. Without loss of generality, let $\mathcal{N}_{\Delta^{\prime}}[G_{1}]=\{G_{1},\ldots,G_{t}\}$ for some $t\leq s$. Since $G_{1}$ is a good leaf, by Proposition˜2.5, we can assume, without loss of generality, that $G_{1}\cap G_{2}\supseteq G_{1}\cap G_{3}\supseteq\cdots\supseteq G_{1}\cap G_{t}$. Now, take any leaf $F$ of $\Delta$ such that $F\cap G_{1}\neq\emptyset$. We first observe that $\mathcal{N}_{\Delta}(F)\subseteq\mathcal{N}_{\Delta^{\prime}}[G_{1}]$. Indeed, if $F\cap G_{j}\neq\emptyset$ for some $j\in[s]\setminus\{1\}$, then since $F$ is a good leaf of $\Delta$, we have either $F\cap G_{1}\supseteq F\cap G_{j}\neq\emptyset$, or $F\cap G_{j}\supseteq F\cap G_{1}\neq\emptyset$. Thus in any case, $G_{1}\cap G_{j}\neq\emptyset$, which in turn, shows that $G_{j}\in\mathcal{N}_{\Delta^{\prime}}[G_{1}]$.
First, we consider the case when $F\nsupseteq G_{1}\cap G_{t}$. We claim that, in this case, $F$ is a special leaf. Note that, by Remark˜2.9, it is enough to show that if $G_{l}\cap G_{m}\neq\emptyset$ for some distinct $l,m\in[s]$, then $(G_{l}\cap G_{m})\setminus F\neq\emptyset$.
Case-I: Either $G_{l}\notin\mathcal{N}_{\Delta^{\prime}}[G_{1}]$, or $G_{m}\notin\mathcal{N}_{\Delta^{\prime}}[G_{1}]$. Without loss of generality, assume $G_{l}\notin\mathcal{N}_{\Delta^{\prime}}[G_{1}]$. Then we have $G_{l}\cap F=\emptyset$ since $\mathcal{N}_{\Delta}(F)\subseteq\mathcal{N}_{\Delta^{\prime}}[G_{1}]$. Hence, $(G_{l}\cap G_{m})\cap F=\emptyset$, which in turn, implies that $(G_{l}\cap G_{m})\setminus F\neq\emptyset$, where $i\in[s]$.
Case-II: $G_{l},G_{m}\in\mathcal{N}_{\Delta^{\prime}}[G_{1}]$. Then, since $G_{1}$ is a good leaf in $\Delta^{\prime}$, without loss of generality, we can assume that $G_{1}\cap G_{l}\supseteq G_{1}\cap G_{m}\supseteq G_{1}\cap G_{t}$. Since $F\nsupseteq G_{1}\cap G_{t}$, there is some $x\in G_{1}\cap G_{t}$ such that $x\notin F$. Observe that $x\in G_{l}\cap G_{m}$, and $x\notin F$. This shows that $x\in(G_{l}\cap G_{m})\setminus F$, where $i\in[s]$. This completes the proof of the claim.
Finally, if $F\supseteq G_{1}\cap G_{t}$, then take any other leaf $F^{\prime}\neq F$ of $\Delta$ such that $F^{\prime}\cap G_{1}\neq\emptyset$. Such a leaf $F^{\prime}$ exists since $G_{1}$ is a facet of $\Delta$. In this case, proceeding as before, we have $\mathcal{N}_{\Delta}(F^{\prime})\subseteq\mathcal{N}_{\Delta^{\prime}}[G_{1}]$. Now since $F\cap F^{\prime}=\emptyset$, we must have $F^{\prime}\nsupseteq G_{1}\cap G_{t}$. Then, we can proceed in a similar way as in the previous case to show that $F^{\prime}$ is a special leaf of $\Delta$. Thus, $\Delta$ contains at least one special leaf. ∎
3. Square-free Powers of Cohen-Macaulay Simplicial Forests
In this section, we explicitly derive the dimension and depth of square-free powers of the facet ideal of a Cohen-Macaulay simplicial forest. As a consequence, we get our main result, which states that all the square-free powers of a Cohen-Macaulay simplicial forest are also Cohen-Macaulay. Additionally, we prove that the normalized depth function of such an ideal is nonincreasing.
Before going to the main content of this section, we clarify some notations to avoid any confusion for the reader. For a square-free monomial ideal $I$, by $I^{[k]}$ we mean the $k^{\text{th}}$ square-free power of $I$, whereas $i\in[k]$ means the element $i$ is in the set $k]:=\{1,2,\ldots,k\}$.
We begin with the following lemma, which shows that for the square-free powers of the facet ideal of a Cohen-Macaulay simplicial forest, taking colon by the monomial associated with a leaf behaves well. Here we assume that $I(\Delta)^{[k]}=R$ for each $k\leq 0$.
Lemma 3.1.
Let $\Delta$ be a Cohen-Macaulay simplicial forest, and let $F$ be a leaf of $\Delta$. Then for all $1\leq k\leq\nu(\Delta)$,
$(I(\Delta)^{[k]}:\mathbf{x}_{F})=I(\Delta_{1})^{[k-1]},$
where $\mathcal{F}(\Delta_{1})=\mathcal{F}(\Delta)\setminus\mathcal{N}_{\Delta}[F]$. Moreover, $\Delta_{1}$ is also a Cohen-Macaulay simplicial forest.
Proof.
The $k=1$ case is easy to observe. Therefore, we may assume that $k\geq 2$. Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$. Without loss of generality, assume that $F=F_{1}$, and $\mathcal{N}_{\Delta}(F_{1})=\{G_{1},\ldots,G_{t}\}$ for some $t\leq s$. Take a minimal generator $\prod_{i=1}^{k}\mathbf{x}_{H_{i}}$ in $I(\Delta)^{[k]}$. Then $\{H_{1},\ldots,H_{k}\}\subseteq\mathcal{F}(\Delta)$ forms a $k$-matching in $\Delta$. If $H_{j}=F_{1}$ for some $j\in[k]$, then $H_{l}\notin\mathcal{N}_{\Delta}(F_{1})$ for each $l\in[k]\setminus\{j\}$. Thus, in this case, we have
$\frac{\prod_{i=1}^{k}\mathbf{x}_{H_{i}}}{\gcd(\mathbf{x}_{F_{1}},\prod_{i=1}^{k}\mathbf{x}_{H_{i}})}=\prod_{\underset{i\neq j}{i=1}}^{k}\mathbf{x}_{H_{i}}\in I(\Delta_{1})^{[k-1]}.$
Thus, we may assume that $H_{j}\neq F_{1}$ for each $j\in[k]$. Now suppose $H_{j}\notin\mathcal{N}_{\Delta}(F_{1})$ for each $j\in[k]$. Then, it is easy to see that
$\frac{\prod_{i=1}^{k}\mathbf{x}_{H_{i}}}{\gcd(\mathbf{x}_{F_{1}},\prod_{i=1}^{k}\mathbf{x}_{H_{i}})}=\prod_{i=1}^{k}\mathbf{x}_{H_{i}}\in I(\Delta_{1})^{[k]}\subseteq I(\Delta_{1})^{[k-1]}.$
Next, consider the case when $H_{j}\in\mathcal{N}_{\Delta}(F_{1})$ for some $j\in[k]$. Observe that if $H_{l}\in\mathcal{N}_{\Delta}(F_{1})$ for some $l\in[k]$ with $l\neq j$, then by Lemma˜2.6, we have either $H_{l}\cap F_{1}\subseteq H_{j}\cap F_{1}$ or $H_{j}\cap F_{1}\subseteq H_{l}\cap F_{1}$. In particular, $H_{l}\cap H_{j}\neq\emptyset$, a contradiction to the fact that $\{H_{1},\ldots,H_{k}\}$ forms a $k$-matching in $\Delta$. Thus $H_{l}\notin\mathcal{N}_{\Delta}[F_{1}]$ for each $l\in[k]\setminus\{j\}$, and hence, $\prod_{\underset{i\neq j}{i=1}}^{k}\mathbf{x}_{H_{i}}\in I(\Delta_{1})^{[k-1]}$. Consequently,
$\frac{\prod_{i=1}^{k}\mathbf{x}_{H_{i}}}{\gcd(\mathbf{x}_{F_{1}},\prod_{i=1}^{k}\mathbf{x}_{H_{i}})}=(\mathbf{x}_{H_{j}\setminus F_{1}})\cdot\left(\prod_{\underset{i\neq j}{i=1}}^{k}\mathbf{x}_{H_{i}}\right)\in I(\Delta_{1})^{[k-1]}.$
Thus considering all the above cases, we have $(I(\Delta)^{[k]}:\mathbf{x}_{F_{1}})\subseteq I(\Delta_{1})^{[k-1]}$. On the other hand, since $\mathcal{F}(\Delta_{1})=\mathcal{F}(\Delta)\setminus\mathcal{N}_{\Delta}[F_{1}]$, we observe that if $H_{1},\ldots,H_{k-1}\in\mathcal{F}(\Delta_{1})$ forms a (k1)(k-1)-matching in $\Delta_{1}$, then $H_{1},\ldots,H_{k-1},F_{1}$ forms a $k$-matching in $\Delta$. Hence, $I(\Delta_{1})^{[k-1]}\subseteq(I(\Delta)^{[k]}:\mathbf{x}_{F_{1}})$. This completes the proof of the first part of the lemma.
It now remains to show that the subcomplex $\Delta_{1}=\langle F_{2},F_{3},\ldots,F_{r}\rangle\cup\langle G_{t+1},G_{t+2},\ldots,G_{s}\rangle$ is also a Cohen-Macaulay simplicial forest. Since $\Delta_{1}$ is a subcomplex of $\Delta$, $\Delta_{1}$ is again a simplicial forest. Thus, it is enough to show that $\Delta_{1}$ is a grafted simplicial complex. By Lemma˜2.6, as $F_{1}$ is a good leaf, without loss of generality, we can assume that $F_{1}\cap G_{1}\supseteq F_{1}\cap G_{2}\supseteq\cdots\supseteq F_{1}\cap G_{t}$. Note that $G_{1}$ is a joint of $F_{1}$, and hence, $\Delta\setminus\langle G_{1}\rangle$ is also grafted. Now, $G_{2}$ is a joint of $F$ in $\Delta\setminus\langle G_{1}\rangle$ and then $\Delta\setminus\langle G_{1},G_{2}\rangle$ is grafted. Continuing in this way, we finally obtain that $\Delta\setminus\langle G_{1},G_{2},\ldots,G_{t}\rangle$ is grafted. Observe that $\Delta\setminus\langle G_{1},G_{2},\ldots,G_{t}\rangle=\langle F_{1}\rangle\cup\Delta_{1}$, where $F_{1}\cap V(\Delta_{1})=\emptyset$. Thus, $\Delta_{1}$ is also grafted, and this completes the proof. ∎
The notion of contraction of a simplicial complex is well-studied in the literature. We recall the definition and prove an auxiliary lemma below that will be essential in the computation of depth.
Definition 3.2 (Contraction).
Let $\Delta$ be a simplicial complex with $\mathcal{F}(\Delta)=\{F_{1},\ldots,F_{l}\}$ and let $A\subseteq V(\Delta)$. The contraction of $\Delta$ on $A$ is a simplicial complex $\Delta_{A}$ on the vertex set $V(\Delta)\setminus A$ whose facets are given by
$\mathcal{F}(\Delta_{A})=\{F_{i}\setminus A\mid F_{i}\in\mathcal{F}(\Delta)\text{ and }F_{i}\setminus A\nsupseteq F_{j}\setminus A\text{ for each }j\neq i\}.$
The following lemma is needed in the proof of Lemma˜3.4.
Lemma 3.3.
Let $I$ be a monomial ideal in a polynomial ring $R$ and let $m\notin I$ be a monomial in $R$. If $R/I$ is Cohen-Macaulay, then so is $R/(I:m)$.
Proof.
By [17, Corollary 1.3] we have $\mathrm{depth}(R/I)\leq\mathrm{depth}(R/(I:m))$. On the other hand, $\mathrm{depth}(R/(I:m))\leq\dim(R/(I:m))\leq\dim(R/I)=\mathrm{depth}(R/I)$ since $R/I$ is Cohen-Macaulay. Thus $\mathrm{depth}(R/(I:m))=\dim(R/(I:m))$ and hence $R/(I:m)$ is Cohen-Macaulay. ∎
Given a Cohen-Macaulay simplicial forest $\Delta$, one can derive that $\Delta_{A}$ is also a Cohen-Macaulay simplicial forest. However, in order to determine the grafting structure in $\Delta_{A}$, one needs to carefully use the grafting structure in $\Delta$. In the next lemma, we have explicitly determined the grafting structure of a special type of contraction in a Cohen-Macaulay simplicial forest that will be essential in the computation of depth in Theorem˜3.6.
Lemma 3.4.
Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ be a Cohen-Macaulay simplicial forest. Take $A_{F_{i}}\subseteq\bigcap_{F_{i}\cap G_{j}\neq\emptyset}(F_{i}\cap G_{j})$. Then $\Delta_{A_{F_{i}}}$ is also a Cohen-Macaulay simplicial forest. Moreover, the structure of $\Delta_{A_{F_{i}}}$ can be explicitly determined.
Proof.
Note that $A_{F_{i}}$ is not a facet of $\Delta$ and thus $\mathbf{x}_{A_{F_{i}}}\notin I(\Delta)$. Since $R/I(\Delta)$ is Cohen-Macaulay, by Lemma˜3.3, we have that $R/(I(\Delta):\mathbf{x}_{A_{F_{i}}})$ is also Cohen-Macaulay. Observe that $I(\Delta_{A_{F_{i}}})=(I(\Delta):\mathbf{x}_{A_{F_{i}}})$. Thus $R/I(\Delta_{A_{F_{i}}})$ is Cohen-Macaulay. We now proceed to explicitly determine the structure of $\Delta_{A_{F_{i}}}$ and show that $\Delta_{A_{F_{i}}}$ is a grafting of some simplicial forest.
Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ with $\Delta^{\prime}=\langle G_{1},\ldots,G_{s}\rangle$ and without loss of generality, we take $F_{i}=F_{1}$. If $A_{F_{1}}=\emptyset$, then $\Delta_{A_{F_{1}}}=\Delta$ and thus $\Delta_{A_{F_{1}}}$ is a grafting of $\Delta^{\prime}$. Therefore, we may assume that $A_{F_{1}}\neq\emptyset$. For simplicity of notation, let us write $A=A_{F_{1}}$. We now describe a procedure to find a minimal presentation of $\Delta_{A}$ as follows.
  • $\bullet$
    Without loss of generality, let us assume that $\mathcal{N}_{\Delta}(F_{1})=\{G_{1},\ldots,G_{n}\}$ for some $n\leq s$. Observe that $G_{l}\setminus A=G_{l}$ for each $l>n$, and $F_{j}\setminus A=F_{j}$ for each $j\neq 1$. Moreover, for each distinct $i,j\in[n]$, $G_{i}\setminus A\nsupseteq G_{j}\setminus A$. Furthermore, for any $i\in[n]$, $F_{j}\setminus A\nsubseteq G_{i}\setminus A$ and $G_{l}\setminus A\nsubseteq G_{i}\setminus A$, where $j\in[r]$ and $l>n$. Thus, $G_{i}\setminus A\in\mathcal{F}(\Delta_{A})$ for each $i\in[n]$.
  • $\bullet$
    Without loss of generality, let for each $n+1\leq l\leq m$, $G_{l}=G_{l}\setminus A\supseteq G_{i}\setminus A$ for some $i\in[n]$; and for each $l>m$ and $i\in[n]$, $G_{l}=G_{l}\setminus A\nsupseteq G_{i}\setminus A$, where $m>n$ is an integer. In this case, we have $G_{l}=G_{l}\setminus A\notin\mathcal{F}(\Delta_{A})$ for each $n+1\leq l\leq m$.
  • $\bullet$
    Since $G_{l}\setminus A\nsupseteq F_{i}\setminus A$ for each $i\in[r]$, we see that $G_{l}\setminus A\in\mathcal{F}(\Delta_{A})$ for each $l>m$.
  • $\bullet$
    Observe that $F_{1}\setminus A\nsupseteq G_{i}\setminus A$ for every $i\in[s]$, and hence, $F_{1}\setminus A\in\mathcal{F}(\Delta_{A})$.
  • $\bullet$
    Without loss of generality, let $2\leq q\leq r$ be an integer such that for each $2\leq t\leq q$, $F_{t}\setminus A\supseteq G_{i}\setminus A$ for some $i\in[n]$; and for each $t>q$ and $i\in[n]$, $F_{t}\setminus A\nsupseteq G_{i}\setminus A$. In this case, $F_{t}\setminus A\notin\mathcal{F}(\Delta_{A})$ for each $2\leq t\leq q$.
  • $\bullet$
    Moreover, since for each $j\in[r]$, $F_{j}\setminus A\nsupseteq G_{l}\setminus A=G_{l}$ for each $l>m$, we have that $F_{t}=F_{t}\setminus A\in\mathcal{F}(\Delta_{A})$ for each $t>q$.
Next, we proceed to show that for each $2\leq t\leq q$, there exists a unique $i_{t}\in[n]$ such that $F_{t}\setminus A\supseteq G_{i_{t}}\setminus A$.
  • $\bullet$
    Observe that $F_{t}\setminus A=F_{t}$ and if $F_{t}\supseteq G_{i_{t}}\setminus A$ and $F_{t}\supseteq G_{j_{t}}\setminus A$ for two distinct $i_{t},j_{t}\in[n]$, then $F_{t}\cap G_{i_{t}}=G_{i_{t}}\setminus A$ and $F_{t}\cap G_{j_{t}}=G_{j_{t}}\setminus A$. Hence, by Lemma˜2.6, either $G_{i_{t}}\setminus A\subseteq G_{j_{t}}\setminus A$ or $G_{i_{t}}\setminus A\supseteq G_{j_{t}}\setminus A$. Thus, either $G_{i_{t}}\subseteq G_{j_{t}}$ or $G_{i_{t}}\supseteq G_{j_{t}}$, a contradiction to the fact that $G_{i}\in\mathcal{F}(\Delta)$ for each $i\in[n]$. Also, since $F_{i}\cap F_{j}=\emptyset$ for each $i\neq j$, we see that if for some $p\in[n]$, $G_{p}\setminus A\subseteq F_{i}$ and $G_{p}\setminus A\subseteq F_{j}$, then we must have $i=j$.
Without loss of generality, we may assume that $F_{t}\setminus A\supseteq G_{t-1}\setminus A$ for each $2\leq t\leq q$. Then there exists non-negative integers $q,n$ and $m$ such that $q\leq\min\{r-1,n\}$, $n\leq s$, $n<m\leq s$, and the minimal presentation of $\Delta_{A}$ is
$\Delta_{A}=\langle F_{1}\setminus A,G_{1}\setminus A,\ldots,G_{q-1}\setminus A,F_{q+1},\ldots,F_{r},G_{q}\setminus A,\ldots,G_{n}\setminus A,G_{m+1},\ldots,G_{s}\rangle.$ (1)
Here we remark that if $A\subsetneq\bigcap_{F_{1}\cap G_{j}\neq\emptyset}(F_{1}\cap G_{j})$, then no such $q$ exists. For simplicity of notation, let us write $F^{\prime}_{i}=F_{i}\setminus A$ and $G^{\prime}_{i}=G_{i}\setminus A$ for all $i$. We now proceed to show that $\Delta_{A}$ is a grafted simplicial forest.
Claim 1: The simplicial complex $\Delta^{\prime}_{A}=\langle G^{\prime}_{q},\ldots,G^{\prime}_{n},G^{\prime}_{m+1},\ldots,G^{\prime}_{s}\rangle$ is a simplicial forest.
Proof of Claim 1: Let $\{G^{\prime}_{j_{1}},\ldots,G^{\prime}_{j_{t}}\}\subseteq\{G^{\prime}_{q},\ldots,G^{\prime}_{n},G^{\prime}_{m+1},\ldots,G^{\prime}_{s}\}$ be an arbitrary collection of facets of $\Delta^{\prime}_{A}$. We have to show the subcomplex $\langle G^{\prime}_{j_{1}},\ldots,G^{\prime}_{j_{t}}\rangle$ of $\Delta^{\prime}_{A}$ has a leaf. Consider the simplicial complex $\langle G_{j_{1}},\ldots,G_{j_{t}}\rangle$, which is a subcomplex of $\Delta^{\prime}$. Since $\Delta^{\prime}$ is a simplicial forest, the subcomplex $\langle G_{j_{1}},\ldots,G_{j_{t}}\rangle$ has a leaf, say $G_{j_{1}}$. Then there exists $i\in\{2\ldots,t\}$ such that $G_{j_{1}}\cap G_{j_{i}}\supseteq G_{j_{1}}\cap G_{j_{k}}$ for all $k\in\{2,\ldots,t\}\setminus\{i\}$. This implies $G^{\prime}_{j_{1}}\cap G^{\prime}_{j_{i}}\supseteq G^{\prime}_{j_{1}}\cap G^{\prime}_{j_{k}}$ for all $k\in\{2,\ldots,t\}\setminus\{i\}$. Hence, $G^{\prime}_{j_{1}}$ is a leaf of the subcomplex $\langle G^{\prime}_{j_{1}},\ldots,G^{\prime}_{j_{t}}\rangle$ of $\Delta^{\prime}_{A}$, and consequently, $\Delta^{\prime}_{A}$ is a simplicial forest.
Claim 2: $\Delta_{A}$ is a grafting of $\Delta^{\prime}_{A}$ with the simplices $F^{\prime}_{1},G^{\prime}_{1},\ldots,G^{\prime}_{q-1},F^{\prime}_{q+1},\ldots,F^{\prime}_{r}$.
Proof of Claim 2: In order to prove Claim 2, we verify conditions (1)-(5) in Definition˜2.1. Since for any two leaves $F_{i}$ and $F_{j}$ of $\Delta$, $F_{i}\cap F_{j}=\emptyset$, we observe that for each $a\in\{1,q+1,\ldots,r\}$ and $b\in\{1,\ldots,q-1\}$, $F_{a}^{\prime}\cap G_{b}^{\prime}=\emptyset$. Moreover, it is easy to see that for each $i,j\in\{1,q+1,\ldots,r\}$, $F_{i}^{\prime}\cap F_{j}^{\prime}=\emptyset$. Furthermore, for each $a,b\in[q-1]$, $G_{a}^{\prime}\cap G_{b}^{\prime}=\emptyset$ since $G_{a}^{\prime}\subseteq F_{a+1}$ and $G_{b}^{\prime}\subseteq F_{b+1}$. Thus, the condition (4) in Definition˜2.1 is verified. Next, for each $i\in\{1,q+1,\ldots,r\}$, using Lemma˜2.6, we obtain a chain of subsets
$F_{i}\cap G_{i_{1}}\supseteq F_{i}\cap G_{i_{2}}\supseteq\cdots\supseteq F_{i}\cap G_{i_{l}},$
where $\mathcal{N}_{\Delta}(F_{i})=\{G_{i_{1}},\ldots,G_{i_{l}}\}$. Consequently, there exists a subset $\{j_{1},\ldots,j_{p}\}\subseteq\{i_{1},\ldots,i_{l}\}$ such that
$F_{i}^{\prime}\cap G_{j_{1}^{\prime}}\supseteq F_{i}^{\prime}\cap G_{j_{2}^{\prime}}\supseteq\cdots\supseteq F_{i}^{\prime}\cap G_{j_{p}}^{\prime},$ (2)
where $\mathcal{N}_{\Delta_{A}}(F_{i}^{\prime})=\{G_{j_{1}}^{\prime},\ldots,G_{j_{p}}^{\prime}\}$. Thus for each $i\in\{1,q+1,\ldots,r\}$, $F_{i}^{\prime}$ is a leaf of $\Delta_{A}$. Now fix some $i\in[q-1]$ so that $F_{i+1}\supseteq G_{i}^{\prime}$. Observe that if $G_{j}^{\prime}\in\mathcal{N}_{\Delta_{A}}(G_{i}^{\prime})$, then $j\in\{q,\ldots,n,m+1,\ldots,s\}$ and $G_{j}\in\mathcal{N}_{\Delta}(F_{i+1})$. By Lemma˜2.6, we again have a chain of subsets
$F_{i+1}\cap G_{l_{1}}\supseteq F_{i+1}\cap G_{l_{2}}\supseteq\dots\supseteq F_{i+1}\cap G_{l_{k}},$
where $\mathcal{N}_{\Delta}(F_{i+1})=\{G_{l_{1}},\ldots,G_{l_{k}}\}$, and $G_{l_{a}}=G_{i}$ for some $a\in[k]$. Consequently, we have the following chain of subsets
$F_{i+1}^{\prime}\cap G_{l_{1}}^{\prime}\supseteq F_{i+1}^{\prime}\cap G_{l_{2}}^{\prime}\supseteq\dots\supseteq F_{i+1}^{\prime}\cap G_{l_{k}}^{\prime}.$ (3)
Note that for each $1\leq b\leq a$, $F_{i+1}\cap G_{l_{b}}\supseteq F_{i+1}\cap G_{l_{a}}=G_{l_{a}}^{\prime}$, and thus $G_{l_{b}}^{\prime}\supseteq G_{l_{a}}^{\prime}$. Consequently, $G_{l_{b}}^{\prime}\notin\mathcal{F}(\Delta_{A})$ if $b<a$. In this case, intersecting $G_{l_{a}}^{\prime}$ with each member of the chain in Equation˜3 we obtain
$G_{l_{a}}^{\prime}\cap G_{l_{a}+1}^{\prime}\supseteq\dots\supseteq G_{l_{a}}^{\prime}\cap G_{l_{k}}^{\prime},$
where $\mathcal{N}_{\Delta_{A}}(G_{l_{a}}^{\prime})\subseteq\{G_{l_{a}+1}^{\prime},\ldots,G_{l_{k}}^{\prime}\}$. Thus we see that $G_{i}^{\prime}$ is a leaf of $\Delta_{A}$ for each $i\in[q-1]$. Therefore, the condition (2) in Definition˜2.1 is verified.
To verify the condition (1) in Definition˜2.1, let us consider a vertex $x\in G_{j}^{\prime}$, where $j\in\{q,\ldots,n,m+1,\ldots,s\}$. Then $x\in G_{j}$. If $x\in F_{i}$ for some $i\in\{1,q+1,\ldots,r\}$, then it is easy to see that $x\in F_{i}^{\prime}$. Now suppose $x\in F_{i}$ for some $i\in\{2,\ldots,q\}$. Note that $F_{i}\cap G_{i-1}\neq\emptyset$, by construction. Thus, in this case, by Lemma˜2.6, either $F_{i}\cap G_{j}\supseteq F_{i}\cap G_{i-1}$ or $F_{i}\cap G_{j}\subseteq F_{i}\cap G_{i-1}$. Observe that if $F_{i}\cap(G_{j}\setminus A)=F_{i}\cap G_{j}\supseteq F_{i}\cap G_{i-1}=G_{i-1}^{\prime}$, then $G_{j}^{\prime}\supseteq G_{i-1}^{\prime}$, a contradiction to the fact that $G_{j}^{\prime}\in\mathcal{F}(\Delta_{A})$. Thus we must have $F_{i}\cap G_{j}\subseteq F_{i}\cap G_{i-1}=G_{i-1}^{\prime}$ and consequently, $x\in G_{i-1}^{\prime}$. Thus, condition (2) in Definition˜2.1 is verified. Also, by the minimal representation of $\Delta_{A}$ in Equation˜1, we see that condition (3) in Definition˜2.1 is also verified.
To prove the condition (5) in Definition˜2.1, first note that if $\Delta^{\prime}$ contains exactly one facet $G_{1}$, then it follows from the definition of grafting that $\Delta_{A}\setminus\langle G_{1}^{\prime}\rangle=(\Delta\setminus\langle G_{1}\rangle)_{A}$ is a grafted simplicial complex. Now suppose $\Delta^{\prime}$ contains more than one facet with $\mathcal{F}(\Delta^{\prime})=\{G_{1},\ldots,G_{s}\}$, as before. Further, suppose for $i\in\{q,\ldots,n\}$, we have $\{G_{i_{1}},\ldots,G_{i_{p}}\}\subseteq\mathcal{F}(\Delta^{\prime})\setminus\{G_{i}\}$ such that $G_{i}\setminus A\subseteq G_{i_{j}}\setminus A$, for each $j\in[p]$. Then for each $i\in\{q,\ldots,n,m+1,\ldots,s\}$, we have
$\Delta_{A}\setminus\langle G_{i}^{\prime}\rangle=\begin{cases}(\Delta\setminus\langle G_{i}\rangle)_{A}&\text{ if }i\in\{m+1,\ldots,s\},\\ (\Delta\setminus\langle G_{i},G_{i_{1}},\ldots,G_{i_{p}}\rangle)_{A}&\text{ if }i\in\{q,\ldots,n\}.\end{cases}$
Here one can see that $A\subseteq\bigcap_{\underset{j\neq i}{F_{1}\cap G_{j}\neq\emptyset}}(F_{1}\cap G_{j})$. Thus, using the fact that $\Delta\setminus\langle G_{i}\rangle$ is grafted for each $i\in[s]$, and by the induction on a number of facets of $\Delta^{\prime}$, we have that $\Delta_{A}\setminus\langle G_{i}^{\prime}\rangle$ is again a grafted simplicial forest. This completes the proof of Claim 2 and, subsequently, the proof of this lemma. ∎
We now proceed to determine the Krull dimension of $R/I(\Delta)^{[k]}$, where $\Delta$ is a Cohen-Macaulay simplicial forest.
Theorem 3.5.
Let $\Delta$ be a Cohen-Macaulay simplicial forest. Then for each $1\leq k\leq\nu(\Delta)$,
$\dim(R/I(\Delta)^{[k]})=|V(\Delta)|-\nu(\Delta)+k-1.$
Proof.
Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ be a Cohen-Macaulay simplicial forest. Then by Remark˜2.7, $\nu(\Delta)=r$. To prove the formula, it is enough to show that $\mathrm{ht}(I(\Delta)^{[k]})=r-k+1$. Let $\mathfrak{p}$ be a minimal prime ideal containing $I(\Delta)^{[k]}$. Recall that, $V(\Delta)\subseteq\cup_{i=1}^{r}F_{i}$. If $\mathrm{ht}(\mathfrak{p})\leq r-k$, then there exist leaves $F_{i_{1}},\ldots,F_{i_{k}}$ of $\Delta$ such that $\mathfrak{p}\cap F_{i_{j}}=\emptyset$ for each $1\leq j\leq k$. In this case $\prod_{j=1}^{k}\mathbf{x}_{F_{i_{j}}}\in I(\Delta)^{[k]}$ but $\prod_{j=1}^{k}\mathbf{x}_{F_{i_{j}}}\notin\mathfrak{p}$, a contradiction. Hence $\mathrm{ht}(\mathfrak{p})\geq r-k+1$ for any minimal prime ideal containing $I(\Delta)^{[k]}$. Now we proceed to construct a minimal prime ideal $\mathfrak{q}$ containing $I(\Delta)^{[k]}$ such that $\mathrm{ht}(\mathfrak{q})=r-k+1$. For each $i\in[r]$, let
[…]
By Lemma˜2.6, we have that $\mathcal{A}_{i}\neq\emptyset$ for each $i\in[r]$. Choose some $x_{i}\in\mathcal{A}_{i}$. Using the above description of the sets $\mathcal{A}_{i}$, we construct the prime ideal
$\mathfrak{q}=\langle x_{i}\in\mathcal{A}_{i}\mid i\in[r-k+1]\rangle.$
Then $\mathrm{ht}(\mathfrak{q})=r-k+1$ and thus it is enough to show that $I(\Delta)^{[k]}\subseteq\mathfrak{q}$. Let $\prod_{i=1}^{k}\mathbf{x}_{H_{i}}\in\mathcal{G}(I(\Delta)^{[k]})$. Then for each $i\in[k]$ there exists a leaf $F_{t_{i}}$ of $\Delta$ such that $H_{i}\cap F_{t_{i}}\neq\emptyset$. We proceed to show that $F_{t_{i}}\neq F_{t_{j}}$ for each $i\neq j$. Observe that if $H_{i}$ is a leaf of $\Delta$ for some $i$, then there is nothing to show since $H_{i},H_{j}$ forms a matching in $\Delta$. Now suppose $H_{i}=G_{t_{i}}$ and $H_{j}=G_{t_{j}}$ so that $G_{t_{i}}\cap F_{t_{i}}\neq\emptyset$ and $G_{t_{j}}\cap F_{t_{j}}\neq\emptyset$. Moreover, we assume $F_{t_{i}}=F_{t_{j}}$. In this case by Lemma˜2.6, $G_{t_{i}}\cap G_{t_{j}}\neq\emptyset$, a contradiction. Thus $F_{t_{i}}\neq F_{t_{j}}$ for each $i\neq j$. It is easy to see that $\mathcal{A}_{t_{i}}\subseteq H_{i}\cap F_{t_{i}}$, and in particular, $x_{i}\in H_{i}\cap F_{t_{i}}$ for each $i\in[k]$. Thus $\prod_{i=1}^{k}\mathbf{x}_{H_{i}}\in\mathfrak{q}$ and this completes the proof. ∎
We are now ready to prove one of the main results of this article, where we establish the combinatorial formula for the depth of the $k^{\text{th}}$ square-free powers of the facet ideal of a Cohen-Macaulay simplicial forest.
Theorem 3.6.
Let $\Delta$ be a Cohen-Macaulay simplicial forest with the vertex set $V(\Delta)$. Let $R=\mathbb{K}[x_{i}\mid x_{i}\in V(\Delta)]$ be the polynomial ring over the field $\mathbb{K}$. Then, for each $1\leq k\leq\nu(\Delta)$,
  1. (i)
    $\mathrm{depth}(R/I(\Delta)^{[k]})=|V(\Delta)|-\nu(\Delta)+k-1,$
  2. (ii)
    $\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F}\rangle))=|V(\Delta)|-\nu(\Delta)+k-1$, where $F$ is a special leaf of $\Delta$.
Proof.
We prove (i) and (ii) simultaneously by induction on $|V(\Delta)|$. If $|V(\Delta)|\leq 2$, then it is easy to verify (i) and (ii). Indeed, if $|V(\Delta)|=1$, then $k=1$ and $I(\Delta)$ is generated by the variable in $R$, and the above inequalities hold. If $|V(\Delta)|=2$, then the only possibilities are $I(\Delta)=\langle x_{1},x_{2}\rangle$ or $I(\Delta)=\langle x_{1}x_{2}\rangle$. When $I(\Delta)=\langle x_{1},x_{2}\rangle$, the possibilities of $k$ are $1$ and $2$, and in case $I(\Delta)=\langle x_{1}x_{2}\rangle$ the only possibility of $k$ is $1$. Thus, one can observe that in all the cases the above equalities hold. Therefore, we may assume that $|V(\Delta)|\geq 3$. In this case, if $k=1$, then $I(\Delta)^{[1]}+\langle\mathbf{x}_{F}\rangle=I(\Delta)$ for any special leaf $F$ of $\Delta$, and thus using Theorem˜3.5, we have
$\mathrm{depth}(R/I(\Delta)^{[1]})$ $=\mathrm{depth}(R/(I(\Delta)^{[1]}+\langle\mathbf{x}_{F}\rangle))$
$=\dim(R/I(\Delta))$
$=|V(\Delta)|-\nu(\Delta).$
In view of this, from now onward, we assume that $k\geq 2$. Let $\Delta=\langle F_{1},\ldots,F_{r}\rangle\cup\langle G_{1},\ldots,G_{s}\rangle$ be a Cohen-Macaulay simplicial forest with $\Delta^{\prime}=\langle G_{1},\ldots,G_{s}\rangle$ and $\Delta$ is a grafting of $\Delta^{\prime}$. Then by Remark˜2.7, $\nu(\Delta)=r$. Now, by Lemma˜2.10, $\Delta$ contains a special leaf. Without loss of generality, let $F_{1}$ be a special leaf of $\Delta$.
Claim : $\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))\geq|V(\Delta)|-r+k-1$.
Proof of Claim: First, we consider the case when $F_{1}\cap G_{i}=\emptyset$ for each $i\in[s]$. In this case, $I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle=I(\Delta_{1})^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle$, where $\mathcal{F}(\Delta_{1})=\mathcal{F}(\Delta)\setminus\{F_{1}\}$. It is easy to see that $\Delta_{1}$ is a Cohen-Macaulay simplicial forest with $|V(\Delta_{1})|=|V(\Delta)|-|F_{1}|<|V(\Delta)|$. Therefore, by the additivity of depth [23, Theorem 3.1.34] and the induction hypothesis, we have
$\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))$ $=\mathrm{depth}(R/(I(\Delta_{1})^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))$
$=\mathrm{depth}(\mathbb{K}[V(\Delta_{1})]/I(\Delta_{1})^{[k]})+\mathrm{depth}(\mathbb{K}[F_{1}]/\langle\mathbf{x}_{F_{1}}\rangle)$
$=|V(\Delta_{1})|-(r-1)+k-1+|F_{1}|-1$
$=|V(\Delta)|-r+k-1.$
Now suppose $F_{1}\cap G_{i}\neq\emptyset$ for some $i\in[s]$. Without loss of generality, let $\mathcal{N}_{\Delta}(F_{1})=\{G_{1},\ldots,G_{n}\}$ and let $A_{F_{1}}=\cap_{F_{1}\cap G_{j}\neq\emptyset}(F_{1}\cap G_{j})$. By Lemma˜2.6, $A_{F_{1}}\neq\emptyset$ and by Lemma˜3.4, $\Delta_{A_{F_{1}}}$ is a Cohen-Macaulay simplicial forest. For the rest of the proof, we follow the notations introduced as in Lemma˜3.4. Thus there exists non-negative integers $q,n$ and $m$ such that $q\leq\min\{r-1,n\}$, $n\leq s$, $n<m\leq s$, and
$\begin{split}\Delta_{A_{F_{1}}}=\langle F_{1}\setminus A_{F_{1}},&G_{1}\setminus A_{F_{1}},\ldots,G_{q-1}\setminus A_{F_{1}},F_{q+1},\ldots,F_{r}\rangle\\ &\langle G_{q}\setminus A_{F_{1}},\ldots,G_{n}\setminus A_{F_{1}},G_{m+1},\ldots,G_{s}\rangle,\end{split}$ (4)
where $\{F_{1}\setminus A_{F_{1}},G_{1}\setminus A_{F_{1}},\ldots,G_{q-1}\setminus A_{F_{1}},F_{q+1},\ldots,F_{r}\}$ are all the leaves of $\Delta_{A_{F_{1}}}$.
Subclaim 1: $(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle):\mathbf{x}_{A_{F_{1}}}=I(\Delta_{A_{F_{1}}})^{[k]}+\langle\mathbf{x}_{F_{1}\setminus A_{F_{1}}}\rangle$.
Proof of Subclaim 1: Let $\prod_{i=1}^{k}\mathbf{x}_{H_{i}}\in\mathcal{G}(I(\Delta)^{[k]})$ such that $\mathbf{x}_{F_{1}}\nmid\prod_{i=1}^{k}\mathbf{x}_{H_{i}}$. Then $\{H_{1},\ldots,H_{k}\}$ is a $k$-matching of $\Delta$ and $F_{1}\neq H_{i}$ for each $i\in[k]$. Note that, if $H_{i}\cap A_{F_{1}}\neq\emptyset$, then $H_{i}\in\mathcal{N}_{\Delta}(F_{1})$. Thus $|\{i\mid H_{i}\cap A_{F_{1}}\neq\emptyset\}|\leq 1$. Hence
[…]
One can observe from Equation˜4 that if $H_{i}\cap A_{F_{1}}=\emptyset$, then either $H_{i}\in\mathcal{F}(\Delta_{A_{F_{1}}})$ or there exists some $H\in\mathcal{F}(\Delta_{A_{F_{1}}})$ such that $H\subseteq H_{i}$. Moreover, if $H_{i}\cap A_{F_{1}}\neq\emptyset$, then $H_{i}\setminus A_{F_{1}}\in\mathcal{F}(\Delta_{A_{F_{1}}})$. Thus using the fact that $\{H_{1},\ldots,H_{k}\}$ forms a $k$-matching in $\Delta$, we obtain $(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle):\mathbf{x}_{A_{F_{1}}}\subseteq I(\Delta_{A_{F_{1}}})^{[k]}+\langle\mathbf{x}_{F_{1}\setminus A}\rangle$.
Let $\prod_{i=1}^{k}\mathbf{x}_{H_{i}\setminus A_{F_{1}}}\in\mathcal{G}(I(\Delta_{A_{F_{1}}})^{[k]})$ such that $\mathbf{x}_{F_{1}\setminus A_{F_{1}}}\nmid\prod_{i=1}^{k}\mathbf{x}_{H_{i}\setminus A}$. In that case $H_{i}\in\mathcal{F}(\Delta)$ and $H_{i}\neq F_{1}$ for each $i\in[k]$. Also, since $F_{1}$ is a special leaf of $\Delta$, we have $H_{i}\cap H_{j}=\emptyset$ for each distinct $i,j\in[k]$. Thus $\{H_{1},\ldots,H_{k}\}$ forms a $k$-matching in $\Delta$. Since $|\{i\mid H_{i}\cap A_{F_{1}}\}|\leq 1$ we see that
[…]
In particular, $\mathbf{x}_{A_{F_{1}}}\cdot\prod_{i=1}^{k}\mathbf{x}_{H_{i}\setminus A_{F_{1}}}\in I(\Delta)^{[k]}$. This completes the proof of Subclaim 1.
Subclaim 2: $F_{1}\setminus A_{F_{1}}$ is a special leaf of the Cohen-Macaulay simplicial forest $\Delta_{A_{F_{1}}}$.
Proof of Subclaim 2: By Lemma˜3.4, $\Delta_{A_{F_{1}}}$ is a Cohen-Macaulay simplicial forest, and $\{F_{1}\setminus A_{F_{1}},G_{1}\setminus A_{F_{1}},\ldots,G_{q-1}\setminus A_{F_{1}},F_{q+1},\ldots,F_{r}\}$ is the set of all the leaves of $\Delta_{A_{F_{1}}}$. By Remark˜2.9, it is enough to show that $(G_{t_{1}}\setminus A_{F_{1}})\cap(G_{t_{2}}\setminus A_{F_{1}})\neq\emptyset$ implies $((G_{t_{1}}\setminus A_{F_{1}})\cap(G_{t_{2}}\setminus A_{F_{1}}))\setminus(F_{1}\setminus A_{F_{1}})\neq\emptyset$, where $t_{1},t_{2}\in\{q,\ldots,n,m+1,\ldots,s\}$ and $t_{1}\neq t_{2}$. Now suppose $(G_{t_{1}}\setminus A_{F_{1}})\cap(G_{t_{2}}\setminus A_{F_{1}})\neq\emptyset$. Hence $G_{t_{1}}\cap G_{t_{2}}\neq\emptyset$. Since $F_{1}$ is a special leaf we have that $(G_{t_{1}}\cap G_{t_{2}})\setminus F_{1}\neq\emptyset$. Moreover, since $A_{F_{1}}\subseteq F_{1}$, we have that $((G_{t_{1}}\cap G_{t_{2}})\setminus F_{1})\setminus A_{F_{1}}=(G_{t_{1}}\cap G_{t_{2}})\setminus F_{1}\neq\emptyset$. In other words, $((G_{t_{1}}\cap G_{t_{2}})\setminus A_{F_{1}})\setminus(F_{1}\setminus A_{F_{1}})\neq\emptyset$. Thus $((G_{t_{1}}\setminus A_{F_{1}})\cap(G_{t_{2}}\setminus A_{F_{1}}))\setminus(F_{1}\setminus A_{F_{1}})\neq\emptyset$. Therefore, we conclude that $F_{1}\setminus A_{F_{1}}$ is a special leaf of $\Delta_{A_{F_{1}}}$.
By the induction hypothesis, the additivity of depth [23, Theorem 3.1.34], Subclaim 1, Subclaim 2, and Lemma˜3.4, we have
$\begin{split}\mathrm{depth}(R/((I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle):\mathbf{x}_{A_{F_{1}}}))&=\mathrm{depth}(\mathbb{K}[V(\Delta)]/(I(\Delta_{A_{F_{1}}})^{[k]}+\langle\mathbf{x}_{F_{1}\setminus A_{F_{1}}}\rangle))\\ &=\mathrm{depth}(\mathbb{K}[V(\Delta_{A_{F_{1}}})]/(I(\Delta_{A_{F_{1}}})^{[k]}+\langle\mathbf{x}_{F_{1}\setminus A_{F_{1}}}\rangle))\\ &\hskip 140.0pt+\mathrm{depth}(\mathbb{K}[A_{F_{1}}])\\ &=|V(\Delta_{A_{F_{1}}})|-r+k-1+|V(\Delta)|-|V(\Delta_{A_{F_{1}}})|\\ &=|V(\Delta)|-r+k-1.\end{split}$ (5)
Next, we have $I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle+\langle\mathbf{x}_{A_{F_{1}}}\rangle=I(\Delta_{2})^{[k]}+\langle\mathbf{x}_{A_{F_{1}}}\rangle$, where $\mathcal{F}(\Delta_{2})=(\mathcal{F}(\Delta)\cup\{A_{F_{1}}\})\setminus\mathcal{N}_{\Delta}[F_{1}]$. Note that for each $H\in\mathcal{F}(\Delta_{2})\cap\mathcal{F}(\Delta)$, we have $H\cap A_{F_{1}}=\emptyset$. Thus, one can see that $A_{F_{1}}$ is a special leaf of $\Delta_{2}$. Moreover, since $\Delta$ is Cohen-Macaulay, $\Delta\setminus\mathcal{N}_{\Delta}[F_{1}]$ is also Cohen-Macaulay. Hence, $\Delta_{2}$ is also a Cohen-Macaulay simplicial forest. Therefore, by the induction hypothesis, we have
$\begin{split}\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle+\langle\mathbf{x}_{A_{F_{1}}}\rangle))&=|V(\Delta_{2})|-r+k-1+|V(\Delta)|-|V(\Delta_{2})|\\ &=|V(\Delta)|-r+k-1.\end{split}$ (6)
Now, we consider the following short exact sequence:
$0\rightarrow R/(J:\mathbf{x}_{A_{F_{1}}})\rightarrow R/J\rightarrow R/(J+\langle\mathbf{x}_{A_{F_{1}}}\rangle)\rightarrow 0,$
where $J=I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle$. Applying the Depth Lemma [23, Lemma 2.3.9], we get
$\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))\geq\min\left\{\mathrm{depth}(R/(J:\mathbf{x}_{A_{F_{1}}})),\mathrm{depth}(R/(J+\langle\mathbf{x}_{A_{F_{1}}}\rangle))\right\}.$
Then by Equation˜5 and (6), we obtain $\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))\geq|V(\Delta)|-r+k-1$. This completes the proof of the Claim.
Again, observe that
$\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))\leq\dim(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))\leq\dim(R/I(\Delta)^{[k]}).$
Thus, by Theorem˜3.5, $\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))\leq|V(\Delta)|-r+k-1$. Consequently,
$\mathrm{depth}(R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle))=|V(\Delta)|-r+k-1.$ (7)
This completes the proof of (ii).
Next, using Lemma˜3.1, we have $(I(\Delta)^{[k]}:\mathbf{x}_{F_{1}})=I(\Delta_{3})^{[k-1]}$, where $\mathcal{F}(\Delta_{3})=\mathcal{F}(\Delta)\setminus\mathcal{N}_{\Delta}[F_{1}]$, and $\Delta_{3}$ is a Cohen-Macaulay simplicial forest. Thus, by the induction hypothesis, $\mathrm{depth}(R/(I(\Delta)^{[k]}:\mathbf{x}_{F_{1}}))=|V(\Delta_{3})|-(r-1)+(k-1)-1+|F_{1}|$. Note that $|V(\Delta_{3})|=|V(\Delta)|-|F_{1}|$. Therefore, we have
$\mathrm{depth}(R/(I(\Delta)^{[k]}:\mathbf{x}_{F_{1}}))=|V(\Delta)|-r+k-1.$ (8)
Now, let us consider the following short exact sequence:
$0\rightarrow R/(I(\Delta)^{[k]}:\mathbf{x}_{F_{1}})\rightarrow R/I(\Delta)^{[k]}\rightarrow R/(I(\Delta)^{[k]}+\langle\mathbf{x}_{F_{1}}\rangle)\rightarrow 0.$
Again, using the Depth Lemma, Equation˜7 and (8), we get $\mathrm{depth}(R/I(\Delta)^{[k]})\geq|V(\Delta)|-r+k-1$. Finally, since $\mathrm{depth}(R/I(\Delta)^{[k]})\leq\dim(R/I(\Delta)^{[k]})$, we conclude in view of Theorem˜3.5 that $\mathrm{depth}(R/I(\Delta)^{[k]})=|V(\Delta)|-r+k-1=|V(\Delta)|-\nu(\Delta)+k-1$. ∎
As an immediate consequence of Theorem˜3.5 and Theorem˜3.6, we achieve our primary goal of the paper:
Corollary 3.7.
Let $\Delta$ be a Cohen-Macaulay simplicial forest. Then for all $k\geq 1$, $R/I(\Delta)^{[k]}$ is also Cohen-Macaulay.
Proof.
If $1\leq k\leq\nu(\Delta)$, then by Theorem˜3.5 and Theorem˜3.6, we have that $R/I(\Delta)^{[k]}$ is Cohen-Macaulay. If $k>\nu(\Delta)$, then $I(\Delta)^{[k]}$ is the zero ideal, and hence, $R/I(\Delta)^{[k]}=R$ is Cohen-Macaulay. ∎
Remark 3.8.
The result of the above corollary does not hold in general for the facet ideals of Cohen-Macaulay simplicial complexes. As simple graphs can be realized as $1$-dimensional simplicial complexes, consider the edge ideal of the graph $G_{1}$ given in Figure˜2. It follows from [22] that $R/I(G_{1})$ is Cohen-Macaulay. On the other hand, $I(G_{1})^{[2]}$ is generated by the monomials $\{x_{1}y_{1}x_{3}y_{3},x_{1}y_{1}x_{2}x_{3},x_{1}y_{1}x_{2}y_{2},x_{3}y_{3}x_{2}y_{2},x_{3}y_{3}x_{1}x_{2},x_{2}y_{2}x_{1}x_{3}\}$. Observe that both $\langle x_{1},x_{2}\rangle$ and $\langle y_{1},y_{2},y_{3}\rangle$ are two minimal prime ideals containing $I(G_{1})^{[2]}$. Thus $I(G_{1})^{[2]}$ is not unmixed and consequently, $R/I(G_{1})^{[2]}$ is not Cohen-Macaulay.
$x_{1}$$x_{2}$$x_{3}$$y_{1}$$y_{2}$$y_{3}$$G_{1}$
$x_{1}$$x_{2}$$x_{3}$$y_{1}$$y_{2}$$y_{3}$$G_{2}$
Figure 2.
Remark 3.9.
For any simplicial complex $\Delta$ the square-free power $I(\Delta)^{[k]}$ is again a square-free monomial ideal, and thus one can associate a simplicial complex $\Delta^{\prime}$ to the ideal $I(\Delta)^{[k]}$. It is interesting to note that although all square-free powers of a Cohen-Macaulay simplicial forest $\Delta$ is again Cohen-Macaulay (by Corollary˜3.7), it may happen that $I(\Delta)^{[k]}$ is not a simplicial forest for some $k$. For instance, consider the edge ideal $I(G_{2})=\langle x_{1}x_{2},x_{2}x_{3},x_{1}y_{1},x_{2}y_{2},x_{3}y_{3}\rangle$ of the graph $G_{2}$ as in Figure˜2. Observe that $I(G_{2})$ is the edge ideal of a whisker three path and thus Cohen-Macaulay [22]. Note that,
$I(G_{2})^{[2]}=\langle x_{1}y_{1}x_{2}y_{2},x_{1}y_{1}x_{3}y_{3},x_{1}y_{1}x_{2}x_{3},x_{1}x_{2}x_{3}y_{3},x_{2}y_{2}x_{3}y_{3}\rangle.$
If we set $I(\Delta^{\prime})=I(G_{2})^{[2]}$, then it is easy to see that $\Delta^{\prime}$ does not contain any free vertex. Thus $I(G_{2})^{[2]}$ is not a facet ideal of a simplicial forest.
In the rest of this section, we deal with the normalized depth function of square-free monomial ideals. Let $I\subseteq R$ be a square-free monomial ideal, and assume that $R$ is the smallest polynomial ring where $\mathcal{G}(I)\subseteq R$. Let $d_{k}$ denote the minimum degree of a monomial belonging to $\mathcal{G}(I^{[k]})$. The depth function of square-free powers of square-free monomial ideals was first considered in [8] (see also [12]). In [8], the authors showed that for any $k\geq 0$, if $I^{[k]}\neq 0$, then one always have $\mathrm{depth}(R/I^{[k]})\geq d_{k}-1$. Thus, it makes sense to define the normalized depth function of $I$ in the following way:
$g_{I}(k)=\mathrm{depth}(R/I^{[k]})-(d_{k}-1),\text{ for }1\leq k\leq\nu(I).$
It was conjectured in [8] that the normalized depth function $g_{I}(k)$ of any square-free monomial ideal $I$ is a nonincreasing function. However, in [19], Fakhari showed that for the class of monomial ideal $I=\langle x_{1}x_{3}x_{i+4},x_{1}x_{4}x_{5},x_{2}x_{3}x_{4},x_{2}x_{3}x_{6}\mid i\in[n-4]\rangle$, $g_{I}(2)-g_{I}(1)$ can be arbitrarily large. Note that $I$ is not the facet ideal of a simplicial forest. Indeed, it is easy to see that the subcollection $\{\{x_{1},x_{3},x_{5}\},\{x_{1},x_{4},x_{5}\},\{x_{2},x_{3},x_{4}\}\}$ does not contain any leaf. In the following theorem, we show that when $\Delta$ is a Cohen-Macaulay simplicial forest, the normalized depth function $g_{I(\Delta)}(k)$ is nonincreasing.
Theorem 3.10.
Let $\Delta$ be a Cohen-Macaulay simplicial forest. Then, the normalized depth function $g_{I(\Delta)}(k)$ is nonincreasing.
Proof.
Note that $g_{I(\Delta)}(k)-g_{I(\Delta)}(k+1)=\mathrm{depth}(R/I(\Delta)^{[k]})-\mathrm{depth}(R/I(\Delta)^{[k+1]})+d_{k+1}-d_{k}$. Since $d_{k+1}-d_{k}\geq 1$, using Theorem˜3.6 we obtain $g_{I(\Delta)}(k)-g_{I(\Delta)}(k+1)\geq 0$, and thus $g_{I(\Delta)}(k)$ is a nonincreasing function. ∎
Acknowledgements. The authors would like to thank the anonymous referees for their careful reading and numerous valuable suggestions, which have significantly enhanced the clarity and readability of the paper. In particular, Remark 2.7, Remark 3.8, and Remark 3.9 arose due to the suggestions of the referees. They also thank Francesco Navarra and Ayesha Asloob Qureshi for carefully going through the first draft of this paper. Das and Roy are partially supported by a grant from the Infosys Foundation. Saha would like to thank the Infosys Foundation for a grant and the NBHM (India) for a Postdoctoral Fellowship during his stay at the Chennai Mathematical Institute.